OpenStudy (anonymous):

What is the limit as X -> 0 of x^tan(x)?

OpenStudy (ikram002p):

tan 0 =0

OpenStudy (ikram002p):

so what u have ill be 0^0 which is indeterminate .

OpenStudy (anonymous):

I think you would do something like... \[ x^{\tan x} = e^{\tan x \ln x} \]

OpenStudy (anonymous):

Well, \(\tan x \) is undefined.

OpenStudy (anonymous):

\(\tan 0\) I mean

OpenStudy (anonymous):

Or is that \(\cot 0\)? Hold on...

OpenStudy (ikram002p):

ermm its zeron ;_;

OpenStudy (anonymous):

It should be a number as it involves L'Hopital's Rule but I can't seem to get past taking the natural log of both sides.

OpenStudy (anonymous):

Well, \(x = e^{\ln x}\) works for \(x>0\), so it can only help us for the one sided limit.

OpenStudy (aum):

\[ y = x^{tan(x)} \\ \ln(y) = tan(x) * \ln(x) \\ \ln(y) = \frac{\ln(x)}{\cot(x)} \\ \lim_{x \rightarrow 0 } \frac{\ln(x)}{\cot(x)} = \lim_{x \rightarrow 0 }1/x / (-\cot(x)\csc(x)) = -\lim_{x \rightarrow 0 }\frac{\sin^2(x)}{x*\cos(x)} = \\ -\lim_{x \rightarrow 0 } \frac{\sin(x)}{x} * \lim_{x \rightarrow 0 } \tan(x) = -1 * 0 = 0\\ y = e^0 = 1 \]